the lateral area of a pyramid with a square base is 240 sq. ft. its base edges are 12 ft. long. what is the slant height of the pyramid?
Answer by ewatrrr(24785)


Hi
lateral area of a pyramid = area of base + area of faces
(240 - 144)/4 = 24 = (1/2)sh *12
48/12 = sh
sh = 4